The perimeter of a rectangular painting is 290 centimeters. If the length of the painting is 94 centimeters, what is its width?
Answer by geetha_rama(94)

Perimerer of rectangle = 2(l+w)
2*(l+w) = 290 cms
length = 94cms
2*(94+w) = 290
w+94 = 145
w = 145 -94
= 51 cms
Width of the rectangle is 51 cms